OpenStudy (anonymous):

A 6-sided number cube is rolled. Find the probability of each event. Write each answer as a fraction. P(5).

OpenStudy (anonymous):

If it is a fair die, the probability of each side is 1/6. P(rolling a 5) = 1/6 In general: p(x) = 1/6 for any x \[\in\] X X = {1, 2, 3, 4, 5, 6}